40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 18:46:06 2011, rana, Update, IOO, MC Locking not working Untitled.pngUntitled.png
    Reply  Wed May 4 11:48:01 2011, Suresh, Update, IOO, MC Locking not working 
Message ID: 4621     Entry time: Wed May 4 11:48:01 2011     In reply to: 4620
Author: Suresh 
Type: Update 
Category: IOO 
Subject: MC Locking not working 

[Valera, Suresh]

The first time I noticed that the MC was not locking was after I had finished switching the RF source installation.  Before this change the RF modulation frequency (for MC) was 29.485 MHz as read from the Marconi RF Source.  We replaced this with a Wenzel crystal source at 29.491 MHz.  This may have changed the loop gain. 

Today, I changed the MC alignment to optimise the MC lock.  Valera pointed out that this is not a desirable solution since it would shift the beam pointing for all components downstream.  However, since we are not sure what was the last stable configuration, we decided to stay with the current settings for now and see the trends of several parameters which would tell us if something is drifting and causing the autolocker to fail.

The MC Auto locker is now working okay.  However to obtain lock initially we reduced the loop gain by decreasing the VCO gain.  We then increased the gain after the autolocker had locked the MC.

 

 

 

ELOG V3.1.3-